# GST: Precise 3D Human Body from a Single Image with Gaussian Splatting Transformers

### v1

@@ 5 Results
In this section, we discuss the results obtained on four datasets in various evaluation settings.
@@
The primary focus of this work is the ability to infer a precise 3D human body from a single image without explicit 3D supervision. We show quantitative results in Table 2 on RenderPeople [1] and HuMMan datasets [4]. We compute the MPJPE error with respect to the ground truth SMPL joints before and after training. The results show that without explicit 3D supervision, our training improves the quality of the pose estimation from the pretrained HMR2 [18] and TokenHMR [12]. Furthermore, Fig. 3 shows some examples of our predictions in comparison with the HMR2 initialization and the ground truth SMPL pose. Our poses visually appear better aligned to the ground truth, emphasizing the results in Table 2. During training, the shape of the SMPL models also changes, with the 3D human shape becoming thinner. SMPL models human body shape without clothing. Our method decouples the body shape from additional layers, such as clothing. We hypothesize that this leads the model to estimate the underlying body shape of the human, effectively using the offsets to model clothes and other deformations.
@@ 5.2 Novel View Synthesis Results
We evaluate our method in the task of novel view synthesis across 4 datasets and compare the results with SHERF [24]. For a fair comparison with our method, which does not assume ground truth SMPL parameters are available, we evaluate SHERF using the estimated HMR2/TokenHMR pose and shape parameters instead of the ground truth ones. Our results are in Tables 2 and 3. Visual results are shown in Figures 4, 5 and 6. Note that the underlying 3D body is consistent, despite a slight blurriness of the Gaussians, and follows precise 3D geometry.
To showcase the capabilities of GST on a large dataset, we train our GST on multi-view images rendered from the TH21 dataset [63], which contains 2,500 3D scans and shows the results on 200 randomly sampled test scans in Table 4 and Fig.7. It clearly shows less blurriness than the other datasets. For reference, we include Splatter Image [56] in Table 4, where our GST predicts precise 3D body pose and shape in addition to the renderable representation unlike Splatter Image.
Figure 4: Single Image NVS. GST on 4 subjects of HuMMan [4] dataset compared to Ground Truth renderings, and SHERF [24] (after being adapted with HMR2 to work with single image input only). GST depicts the correct human pose (compared with ground truth).
Figure 5: Single Image NVS on 2 subjects of Zju-Mocap [43] compared to SHERF [24] (after being adapted with HMR2 to work with single image input only). GST shows improved visual quality, especially when comparing the depicted pose to ground truth.
Figure 6: Visualization of Single Image Novel View Synthesis Results on RenderPeople. We show single image novel view synthesis results on two subjects of RenderPeople [1] dataset of our GST (top row) compared to Ground Truth renderings (bottom row) of each subject.
@@ 5.3 Ablation and analysis
Input Image GT SMPL Out SMPL SMPL overlays
@@
Ablation Study. We present an ablation study of different design choices and key elements in our architectures and losses and their effect on the 2D and 3D results of a single image to 3D of humans on the HuMMan dataset [4] in Table 5. For these experiments, we report PSNR, SSIM and LPIPS metrics computed on the entire image. It shows the importance of combining the LPIPS, tightness, and transparency loss on the final 3D precision while maintaining the visual fidelity intact. The tightness regularization of (3) has the highest impact on 3D precision, as it favors solutions in which the majority of the pose corrections are obtained with the SMPL parameters, and the Gaussians are only used for small refinements. In contrast, removing the tightness regularization encourages unrealistic and less precise poses, with much larger adjustments obtained with the Gaussian offsets. We also visualize this effect in Fig. 8. We conduct additional ablations in the Appendix.
3D Pose Estimation from Sparse Views. We train GST on the common 3D pose estimation dataset Human3.6M [5] using the default split for train and test subjects (subjects 9 and 11 are used for testing). This dataset is not ideal for our method as it only has 4 views and very few subjects, therefore it’s difficult to generalise to unseen poses and subjects. Additionally, the human masks provided with the dataset are not always precise and our method tends to model parts of the background together with the human. This affects the quality of both the visual results and the 3D pose estimation. The visual metrics for our GST are evaluated on a squared crop of size 256x256 around the human with a PSNR of 18.68 and a 3D error of MPJPE $\downarrow$ = 63.7 mm compared to 50.0 mm for HMR2 [18].
@@ 6 Conclusions and Discussion
In this paper, we introduced GST, a novel approach for human 3D representation that predicts 3D Gaussian Splatting [28], enabling fast rendering with accurate poses. GST leverages multi-view supervision to accurately predict 3D joint and body poses while preserving the perceptual quality of novel view renderings. This dual capability combines precise pose estimation with high-quality rendering, bridging two research paradigms and showcasing the benefits of our approach (See Fig. 3).
Limitations. The main limitation in our method is the requirement of multi-view datasets to train. Another issue is the slight blurriness that appears on some of the renderings as a result of the generalization limitation of the transformer trained on very small datasets we employ (in terms of subject diversity). A possible solution to this is to use multiple datasets or bigger datasets.
@@ A.2 Overfitting Example
To test that the number of Gaussians is sufficient to produce sharp details, we train our model to overfit a single data sample. We obtain an almost perfect reconstruction with PSNR of 41. Image I shows examples of the renderings we obtained. This result confirms our assumption that with a large enough dataset, our model would be able to learn sharper details than it currently learns on the small scale datasets.
@@ A.3 Additional Details for TH21 Experiment
For the TH21 [63] experiment in Table 4 in the main report, we use 72 views rendered in a loop around the subject. We train both our method and Splatter Image [56] using 256x256 images. Despite our model performing worse than Splatter Image in terms of visual metrics, our model also predicts the SMPL paramters for 3D pose estimation. This is both useful for downstream tasks, but also ensures that the underlying 3D shape is plausible for a human. This difference can be noticed in the examples in Figure II, where GST can reconstruct a plausible human shape despite the uncommon input pose, while Splatter Image fails to reconstruct arms and legs.
Refer to captionSplatter ImageGST (ours)Ground truth
@@
A.4 HMR2 Finetuning Comparison
As we discussed in the main paper, we train GST starting from a pretrained version of HMR2 [18] and the resulting model almost halves the MPJPE error on the two datasets, compared to the original pretrained version. We now instead compare GST with a finetuned version of HMR2 to test the quality of the 3D pose estimation of our method.
Table II: HMR2 finetuning comparison MPJPE (in mm) for 3D keypoints evaluation.
Method 3D annotations RenderPeople HuMMan
MPJPE (mm)$\downarrow$ MPJPE (mm)$\downarrow$
GST 64.6 67.6
HMR2 [18] pretrained 101.0 133.4
HMR2 [18] finetune w/ 2D data 127.40 163.77
HMR2 [18] finetune w/ 2D + 3D data 57.33 61.20
We finetune HMR2 on the two datasets in Table 2 in the main paper: HuMMan [4] and RenderPeople [1]. The results are reported in Table II. To reproduce a similar training setup to our method (that does not require any 3D ground truth annotations), we finetune HMR2 using only 2D keypoints annotations. We use images from all views in the dataset, but restrict the supervision to only use the 2D keypoints loss.
The results show that the 2D information alone is not enough for HMR2 to improve the quality of the 3D pose estimation on the two datasets, and the finetuned model MPJPE error is worse than the pretrained one.
For completeness, we also report the errors when finetuning HMR2 with additional 3D annotations: 3D keypoints and ground truth SMPL parameters. We would like to emphasize that we think this is an unfair comparison to our method, since our method does not use ground truth SMPL parameters or 3D keypoints for training. The MPJPE of the HMR2 version finetuned with 3D data is only 7mm better than ours on RenderPeople and 6mm better than ours on HuMMan.
@@ Appendix B Additional Visualizations
We provide some additional examples of novel view synthesis and 3D pose estimation results.
Figure III: Results in TH21 [63]. Rendering results for GST (top row) compared to Ground Truth renderings (bottom row) of each subject. An example of loose clothes is in the last row.
Input Image GT View Out View GT Ours HMR2 [18] SMPL Overlays
Figure IV: Additional 3D SMPL Shape Results for the RenderPeople dataset [1]. We show 3D human body results of our GST on three subjects of RenderPeople [1] dataset compared to Ground Truth renderings, Ground Truth SMPL parameters [37], and SMPL predictions of HMR2 [18]. The overlay of the 3 SMPL bodies ( ours, HMR2, and GT) shows that our predicted SMPL is more precise while our predicted Gaussian splats maintain visual quality from novel views.
Input Image GT View Out View GT Ours HMR2 [18] SMPL Overlays
Figure V: Additional 3D SMPL Shape Results for the HuMMan dataset [4]. We show 3D human body results of our GST on three subjects of HuMMan [4] dataset compared to Ground Truth renderings, Ground Truth SMPL parameters [37], and SMPL predictions of HMR2 [18]. The overlay of the 3 SMPL bodies ( ours, HMR2, and GT) shows that our predicted SMPL is more precise while our predicted Gaussian splats maintain visual quality from novel views.
Figure VI: Visualization of Single Image Novel View Synthesis Results on HuMMan. We show single image novel view synthesis results on 7 subjects of HuMMan [4] dataset of our GST (top row) compared to Ground Truth renderings (bottom row) of each subject.
Figure VII: Visualization of Single Image Novel View Synthesis Results on THuman. We show single image novel view synthesis results on 5 subjects of THuman [72] dataset of our GST (top row) compared to Ground Truth renderings (bottom row) of each subject.
Figure VIII: Visualization of Single Image Novel View Synthesis Results on RenderPeople. We show single image novel view synthesis results on 6 subjects of RenderPeople [1] dataset of our GST (top row) compared to Ground Truth renderings (bottom row) of each subject.

### v2

@@ 5 Results
In this section, we discuss the results obtained on five datasets in various evaluation settings.
@@
The primary focus of this work is the ability to infer a precise 3D human body from a single image without explicit 3D supervision. We show quantitative results in Tab. 2 on RenderPeople [1] and HuMMan datasets [2]. We compute the MPJPE error with respect to the ground truth SMPL joints before and after training. The results show that without explicit 3D supervision, our training improves the quality of the pose estimation from the pretrained HMR2 [12] and TokenHMR [7]. Furthermore, Fig. 3 shows some examples of our predictions in comparison with the HMR2 initialization and the ground truth SMPL pose. Our poses appear visually to be better aligned with the ground truth, highlighting the results in Tab. 2.
Refer to captionRefer to captionInput ImageOther ViewsGTGST (ours)vs GTHMR2-3D finetuningvs GTHMR2-2D finetuningvs GT
Figure 4: 3D Shape Comparison with HMR2 After Fine-tuning on 2D and 3D Data. 3D human body results of our GST on two subjects of HuMMan [2] dataset compared to Ground Truth SMPL parameters [29], and SMPL predictions of HMR2 [12]. We show two versions of HMR2, one finetuned on 2D data only (HMR2-2D), and one finetuned on 3D data (HMR2-3D). Our method is only finetuned on 2D image data, but the results are almost as accurate as HMR2 finetuned on 3D data.
We also compare our results against a fine-tuned version of HMR2 (c.f. Tab. 2). To reproduce a similar training setup to our method (that does not require any 3D ground truth annotations), we fine-tune HMR2 using only 2D keypoint annotations. We use images from all views in the dataset but restrict the supervision only to use the 2D keypoints loss. The results show that the 2D information alone is not enough for HMR2 to improve the quality of the 3D pose estimation on the two datasets, and the fine-tuned model MPJPE error is worse than the pretrained one. For completeness, we also report the errors when fine-tuning HMR2 with additional 3D annotations: 3D keypoints and ground truth SMPL parameters. We would like to emphasize that we think this is an unfair comparison to our method since our method does not use ground truth SMPL parameters or 3D keypoints for training. The MPJPE of the HMR2 version, fine-tuned with 3D data, is only 7mm better than ours on RenderPeople and 6mm better than ours on HuMMan. Figure 4 shows some examples of our predictions compared to the HMR2 models finetuned on 2D and 3D data.
@@ 5.2 Novel View Synthesis Results
We evaluate our method in the task of novel view synthesis across 4 datasets and compare the results with SHERF [15]. For a fair comparison with our method, which does not assume ground truth SMPL parameters are available, we evaluate SHERF using the estimated HMR2/TokenHMR pose and shape parameters instead of the ground truth ones. Our results are in Tables Tabs. 2 and 3. Visual results are shown in Fig. Fig. 5. Note that the underlying 3D body is consistent, despite a slight blurriness due to the Gaussians, and follows precise 3D geometry.
To showcase the capabilities of GST on a large dataset, we train our GST on multi-view images rendered from the TH21 dataset [46], which contains 2,500 3D scans and shows the results on 200 randomly sampled test scans in Tab. 4 and Fig. 6. It clearly shows less blurriness than the other datasets. For reference, we include Splatter Image [41] in Tab. 4, where our GST predicts precise 3D body pose and shape in addition to the renderable representation, unlike Splatter Image. Predicting the human model parameters is not only useful for downstream tasks but also ensures the reconstructed shape is plausible for a human. We include additional details on the TH21 experiment and the comparison with Splatter Image in the Appendix.
Refer to captionRefer to captionInput ImageSHERF [15] w/ [12]GST (ours)GT
Figure 5: Single Image NVS. GST on 2 subjects of HuMMan [2] dataset compared to Ground Truth renderings, and SHERF [15] (after being adapted with HMR2 to work with single image input only). GST depicts the correct human pose (compared with ground truth).
@@ 5.3 Ablation and analysis
Refer to captionInput ImageGT SMPLOur SMPL SMPL overlays
@@
Ablation Study. We present an ablation study of different design choices and key elements in our architectures and losses and their effect on the 2D and 3D results of a single image to 3D of humans on the HuMMan dataset [2] in Tab. 5. For these experiments, we report PSNR, SSIM, and LPIPS metrics computed for the entire image. It shows the importance of combining the LPIPS, tightness, and transparency loss on the final 3D precision while maintaining the visual fidelity intact. The tightness regularization of (3) has the highest impact on 3D precision, as it favors solutions in which the majority of the pose corrections are obtained with the SMPL parameters, and the Gaussians are only used for small refinements. In contrast, removing the tightness regularization encourages unrealistic and less precise poses, with much larger adjustments obtained with the Gaussian offsets. We also visualize this effect in Fig. 7. We conduct additional ablations in the Appendix, and we also report some results for 3D pose estimation from sparse views on the Human3.6M dataset [3].
@@ 6 Conclusions and Discussion
In this paper, we introduced GST, a novel approach for human 3D representation that predicts 3D Gaussian Splatting [22], enabling fast rendering with accurate poses. GST leverages multi-view rendering supervision to refine the 3D joint and body predictions. This dual capability combines precise pose estimation with novel view rendering, bridging two research paradigms and showcasing the benefits of our approach. Our method could be used to reduce the friction of training a pose estimation model to deploy to any novel sports setup, without requiring any expensive pose annotations or 3D scans.
Limitations. The main limitation in our method is the requirement of multi-view datasets to train. Another issue is the slight blurriness that appears in some of the renderings, possibly as a result of the generalization limitations of the transformer when trained on relatively small datasets (in terms of subject diversity). A possible solution to this is to use a larger dataset or to pretrain on synthetic data.
[…]
Supplementary Materials
@@ A.2 Overfitting Example
To test that the number of Gaussians is sufficient to produce sharp details, we train our model to overfit a single data sample. We obtain an almost perfect reconstruction with PSNR of 41. Fig. I shows examples of the renderings we obtained. This result confirms our assumption that with a large enough dataset, our model would be able to learn sharper details than it currently learns on the small scale datasets.
@@ A.3 Additional Details for TH21 Experiment
For the TH21 [46] experiment in Table 4 in the main report, we use 72 views rendered in a loop around the subject. We train both our method and Splatter Image [41] using 256x256 images. Despite our model performing worse than Splatter Image in terms of visual metrics, our model also predicts the SMPL paramters for 3D pose estimation. This is both useful for downstream tasks, but also ensures that the underlying 3D shape is plausible for a human. This difference can be noticed in the examples in Fig. II, where GST can reconstruct a plausible human shape despite the uncommon input pose, while Splatter Image fails to reconstruct arms and legs.
Refer to captionRefer to captionSplatter ImageGST (ours)Ground truth
@@
A.4 3D Pose Estimation from Sparse Views
We train GST on the common 3D pose estimation dataset Human3.6M [3] using the default split for train and test subjects (subjects 9 and 11 are used for testing). This dataset is not ideal for our method as it only has 4 views and very few subjects, therefore it’s difficult to generalize to unseen poses and subjects. Additionally, the human masks provided with the dataset are not always precise and our method tends to model parts of the background together with the human. This affects the visual results and the 3D pose estimation. The visual metrics for our GST are evaluated on a squared crop of size 256x256 around the human with a PSNR of 18.68 and a 3D error of MPJPE $\downarrow$ = 63.7 mm compared to 50.0 mm for HMR2 [12].
Refer to captionInput ImageOther ViewsHMR2 [12]GST (ours)Input ImageOther ViewsHMR2 [12]GST (ours)
Figure III: Example of human pose improvements using our method GST. 3D human body results of our GST and SMPL predictions of HMR2 [12] on a sports sequence from the CMU panoptic dome dataset [20].
@@ Appendix B Additional Visualizations
Fig. III shows additional pose estimation results on the sports sequence of the CMU Panoptic dataset [20]. Fig. IV shows additional examples of novel view synthesis comparisons with SHERF [15]. Fig. V and VI show additional pose comparisons for the RenderPeople [1] dataset. Fig. VII and VIII show examples of novel view synthesis results for the TH21 [46], THuman [55] and RenderPeople [1] datasets.
Figure IV: Single Image NVS on two subjects of Zju-Mocap [32] and two subjects of HuMMan [2] compared to SHERF [15] (after being adapted with HMR2 to work with single image input only). GST shows improved visual quality, especially when comparing the depicted pose to ground truth.
Refer to captionRefer to captionInput ImageOther ViewsGTGST (ours)vs GTHMR2 [12]vs GT
Figure V: 3D Shape Comparison with HMR2. 3D human body results of our GST on two subjects of RenderPeople [1] dataset compared to Ground Truth SMPL parameters [29], and SMPL predictions of HMR2 [12].
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionInput ImageOther ViewsGTGST (ours)vs GTHMR2-3D finetuningvs GTHMR2-2D finetuningvs GT
Figure VI: 3D Shape Comparison with HMR2 After Fine-tuning on 2D and 3D Data. 3D human body results of our GST on five subjects of RenderPeople [1] dataset compared to Ground Truth SMPL parameters [29], and SMPL predictions of HMR2 [12]. We show two versions of HMR2, one finetuned on 2D data only (HMR2-2D), and one finetuned on 3D data (HMR2-3D). Our method is only finetuned on 2D image data, but the results are almost as accurate as HMR2 finetuned on 3D data.
Figure VII: Results in TH21 [46]. Rendering results for GST (top row) compared to Ground Truth renderings (bottom row) of each subject. An example of loose clothes is in the last row.
Figure VIII: Visualization of Single Image Novel View Synthesis Results on THuman and RenderPeople. We show single image novel view synthesis results on one subject of THuman [55] dataset and one subjects of RenderPeople [1] dataset of our GST (top row) compared to Ground Truth renderings (bottom row) of each subject.